# Quantum-optical theory of the few femtosecond nonlinear optical response of Drude metals with a non-parabolic conduction band

### v1

@@ II.3 Calculation of internal fields of a single small nanosphere in the time domain
Now that we have obtained the dynamics of the DM elements, to obtain the electron and electromagnetic field dynamics, one generally requires a self-consistent numerical iteration scheme involving the Boltzmann equation, Maxwell’s equations, and Eq. (14). For simple geometries, it is possible to circumvent the direct solution of Maxwell’s equations by leveraging existing known solutions. As an illustrative example, we consider a single small spherical NP with a radius $a$ much smaller than the central wavelength of the incident pulse $\lambda_{0}$, embedded in a homogeneous host medium of permittivity $\varepsilon_{h}$, see Fig. 2(a). This enables employing the quasi-electrostatic approximation of Maxwell’s equations [58], i.e., $\partial{\bf B}/\partial t\approx 0$. We decompose the total polarization density into non-dispersive and dispersive contributions. The non-dispersive part accounts for the response of the bound electrons and is assumed, due to the interest in near infrared frequencies, to remain unchanged under illumination; it is characterized by a constant susceptibility $\chi_{\infty}=\varepsilon_{\infty}-1$, where $\varepsilon_{\infty}$ is the permittivity at infinite frequency. The dispersive contribution, denoted by ${\bf P}_{\textrm{NP}}$, arises from the conduction electrons and is obtained from Eq. (15). The advantage of this configuration is that the local electric field ${\bf E}_{\textrm{NP}}$, (the dispersive part of) the polarization density ${\bf P}_{\textrm{NP}}$ as well as the electric displacement ${\bf D}_{\textrm{NP}}=(\varepsilon_{0}\varepsilon_{\infty}{\bf E}_{\textrm{NP}}+{\bf P}_{\textrm{NP}})$ inside the NP are uniform and that they relate to the incident electric field via [58]
@@
where $\varepsilon_{0}$ is the vacuum permittivity. In the weak-field and continuous-wave limit, the Fourier transform of ${\bf P}_{\textrm{NP}}$ and Eq. (17) recover, respectively, the Drude susceptibility multiplied by the Fourier-transformed local electric field (see details in Ref. [28]) and the standard quasi-electrostatic solution for the internal field of a small spherical NP [59].
@@
Figure 2(b) shows the ITO permittivity and the corresponding prefactor of the linear quasi-static local-field solution in the weak-field, continuous-wave limit. The electric field of the incident pulse is assumed to have a Gaussian envelope, i.e., ${\bf E}_{\textrm{inc}}(t)=\hat{{\bf z}}E_{0}e^{-2\ln 2(t/\tau_{p})^{2}}\cos(\omega_{0}t)$, where $E_{0}$ is the peak amplitude of the incident pulse. We set the pulse duration to $\tau_{p}=5$ fs and choose a central frequency $\omega_{0}/2\pi=175$ THz (equivalently $\sim 0.723$ eV), close to the resonance (see Fig. 2(b)), corresponding to an optical cycle of approximately 6 femtoseconds. The incident pulse is assumed to be polarized in the $z$ direction without loss of generality. However, for simplicity, from now on we suppress the vectorial nature of the electric fields.
@@ III.1 Term comparison
[…]
Figure 3: (a)-(c) The photon-e term (blue), e-e (orange), and e-ph (green) terms for $E_{0}=0.6$ V/nm at t=3.2t=-3.2 fs, 0.5-0.5 fs, and $2.2$ fs. (d) The color map of the photon-e term (Eq. (13)). The black dashed lines label the time sections of (a)-(c). The blue and orange dashed lines label the energy sections shown in (e). (e) The electron distribution dynamics at $\mathcal{E}=\mathcal{E}_{F}\pm\hbar\omega_{0}/8$ (blue and orange).
@@
Thus, specifically, Figs. 3(d) and (e) reveal time-oscillations at frequency $\sim 2\omega_{0}$ (i.e., with a period of $\sim 3$ fs) in this term. These arise from the product between the $\pm\omega_{0}$ components of the real electric field and the DM off-diagonal elements, which themselves oscillate at $\pm\omega_{0}$ as driven by the field (see Eqs. (13) and (14)). This $\sim 2\omega_{0}$ oscillation in the photon-e term, in turn, induces similar oscillations in the distribution. Simultaneously, cross-terms between the $\pm\omega$ components of $E(t)$ and the $\mp\omega$ components of the DM off-diagonal elements generate a zero-frequency component in the photon-e term, resulting in a non-vanishing, time-averaged change in the distribution. As the field intensity increases, the electron distribution oscillation stimulates higher-order oscillations in the off-diagonal density matrix at $\sim\pm 3\omega_{0}$, $\sim\pm 5\omega_{0}$, etc., via Eq. (10b). This cascade gives rise to additional photon-e and distribution oscillations at $4\omega_{0}$, $6\omega_{0}$, and beyond, see Fig. 4. These high-frequency oscillations are typically neglected in the rotating wave approximation; however, in the current context, they signify the crucial role played by the off-diagonal elements of the density matrix in the time evolution. Notably, the off-diagonal elements are mostly dominated by those near the diagonal, with magnitudes comparable to the diagonal elements. More specifically, only off-diagonal elements near the Fermi energy are non-zero at the leading edge of the pulse, while elements further below the Fermi energy become pronounced towards the trailing edge of the pulse, see Fig. 5.
@@ III.2 Excited state absorption
$f(\mathcal{E},t)-f(\mathcal{E}^{\prime},t)=f(\mathcal{E},t)(1-f(\mathcal{E}^{\prime},t))\Theta(\mathcal{E}^{\prime}-\mathcal{E})+f(\mathcal{E},t)(1-f(\mathcal{E}^{\prime},t))\Theta(\mathcal{E}-\mathcal{E}^{\prime})\\ +(1-f(\mathcal{E},t))f(\mathcal{E}^{\prime},t)\Theta(\mathcal{E}-\mathcal{E}^{\prime})+(1-f(\mathcal{E},t))f(\mathcal{E}^{\prime},t)\Theta(\mathcal{E}^{\prime}-\mathcal{E}).$ (20)
If we then multiply each of these four terms by the electron energy and the eDOS, and then integrate them over time (rather than on electron energy as in Eq. (18)), the resulting quantities represent the individual contributions of each process to the total energy change of the electron system, i.e.,
@@
i.e., the sum of all four contributions, $u_{\textrm{abs,in}}$, $u_{\textrm{abs,out}}$, $u_{\textrm{em,in}}$, and $u_{\textrm{em,out}}$, integrated over $\mathcal{E}$ recovers the total energy transferred via the photon–e interaction, i.e., Eq. (19). Figs. 7(b) and (c) show the contributions from absorption (out) and emission (out) for $E_{0}=0.6$ V/nm and $1.2$ V/nm.Figs. 7(b)-(e) show the contributions from absorption (in and out) and emission (in and out) for $E_{0}=0.6$ V/nm and $1.2$ V/nm. At the lower illumination level ($E_{0}=0.6$ V/nm), absorption is dominated by excitation from below the Fermi level to higher-energy states. As the field strength increases, this contribution weakens, and excitations from already excited electrons ($\mathcal{E}>\mathcal{E}_{F}$) to even higher levels become more significant, see Fig. 7(b). The corresponding absorption (in) process exhibits behavior similar to that of the absorption (out) counterparts. In particular, the energy-resolved contribution $u_{\textrm{abs (in)}}(\mathcal{E})$ peaks near $\mathcal{E}_{F}+\hbar\omega_{0}$. As the illumination level increases, the peak slightly decreases, while the contributions at higher energies ($\mathcal{E}\gtrsim\mathcal{E}_{F}+\hbar\omega_{0}$) become more prominent, again indicating the onset of excited state absorption processes. This behavior indicates that the superlinear scaling of (stimulated) absorption with incident power originates from a non-simultaneous two-photon absorption process, analogous to excited-state absorption in atomic systems. It reflects the increasing availability of high-energy absorption channels at elevated field strengths and contrasts sharply with the saturation behavior typical of two-level systems, where absorption eventually becomes sublinear due to ground-state depletion.
In contrast, the emission (out) (and (in)) contributions, when normalized by $|E_{0}|^{2}$, increase with incident power, but their energy dependence remains largely unchanged, see Fig. 7(c). This is consistent with the observed $|E_{0}|^{4}$ scaling of emission shown in Fig. 4(e) of Ref. [1]. Notably, such behavior can only be captured by explicitly tracking the time-dependent population dynamics and cannot be reproduced by phenomenological models.
@@ III.3 Local field
Figure 9: (a) The phase of the prefactor of the linear quasi-static solution for the local field inside a subwavelength dielectric sphere; the vertical dashed line represents the central frequency ($175$ THz) used in the simulations. (b) The phase difference between the local field $E_{\textrm{NP}}(t)$ (17) and the incident field for the three cases shown in Fig. 8. (c) The phase diference between the local and incident fields under weak illumination (i.e., in the linear regime) for three pulse durations: 5 fs (solid), 20 fs (dashed), and 80 fs (dash-dotted).
@@ IV Outlook
This work paves the way to more comprehensive studies of the nonlinear optical response for different particle sizes, different material parameters or TCOs (effective mass, non-parabolicity, electron density, …) different geometries (primarily, of layers, which is the geometry of choice in most experiments to date), for different pump wavelengths (e.g., for a $800$ nm, as in [19]), probe wavelength and their duration and chirp. It would also be a basis for comparison with different more computationally accessible approaches (e.g., the hydrodynamic approach, see [69]), thus, enabling further studies of complex spatio-temporal dynamics of single cycle pulses in TCOs and interplay of coherent wave mixing with permittivity changes studied therein.

### v2

@@ II.3 Calculation of internal fields of a single small nanosphere in the time domain
Now that we have obtained the dynamics of the DM elements, to obtain the electron and electromagnetic field dynamics, one generally requires a self-consistent numerical iteration scheme involving the Boltzmann equation, Maxwell’s equations, and Eq. (14). For simple geometries, it is possible to circumvent the direct solution of Maxwell’s equations by leveraging existing known solutions. As an illustrative example, we consider a single small spherical NP with a radius $a$ much smaller than the central wavelength of the incident pulse $\lambda_{0}$, embedded in a homogeneous host medium of permittivity $\varepsilon_{h}$, see Fig. 2(a). This enables employing the quasi-electrostatic approximation of Maxwell’s equations [63], i.e., $\partial{\bf B}/\partial t\approx 0$. We decompose the total polarization density into non-dispersive and dispersive contributions. The non-dispersive part accounts for the response of the bound electrons and is assumed, due to the interest in near infrared frequencies, to remain unchanged under illumination; it is characterized by a constant susceptibility $\varepsilon_{\infty}-1$, where $\varepsilon_{\infty}$ is the permittivity at infinite frequency. The dispersive contribution, denoted by ${\bf P}_{\textrm{NP}}$, arises from the conduction electrons and is obtained from Eq. (15). The advantage of this configuration is that the local electric field ${\bf E}_{\textrm{NP}}$, (the dispersive part of) the polarization density ${\bf P}_{\textrm{NP}}$ as well as the electric displacement ${\bf D}_{\textrm{NP}}=(\varepsilon_{0}\varepsilon_{\infty}{\bf E}_{\textrm{NP}}+{\bf P}_{\textrm{NP}})$ inside the NP are uniform and that they relate to the incident electric field via [63]
@@
where $\varepsilon_{0}$ is the vacuum permittivity. In the weak-field and continuous-wave limit, the Fourier transform of ${\bf P}_{\textrm{NP}}$ (Eq. (15)) recovers the Drude susceptibility multiplied by the Fourier-transformed local electric field. The resulting permittivity can be written as  [^10] In Eq. (2), the local electric field is assumed in $z$-direction. In this regard, Eq. (18) represents the $zz$-component of the permittivity tensor. (see also details in Ref. [28])
$\varepsilon(\omega,T_{0})=\varepsilon_{\infty}+\dfrac{V}{2\varepsilon_{0}\hbar}\int d\mathcal{E}d\mathcal{E}^{\prime}\dfrac{\rho_{e}(\mathcal{E})\rho_{e}(\mathcal{E}^{\prime})(\mathcal{E}-\mathcal{E}^{\prime})|D(k(\mathcal{E}),k(\mathcal{E}^{\prime}),L)|^{2}(f(\mathcal{E},T_{0})-f(\mathcal{E}^{\prime},T_{0}))}{\omega^{2}-(\mathcal{E}-\mathcal{E}^{\prime})^{2}/\hbar^{2}-2i\eta_{\mathcal{E}\mathcal{E}^{\prime}}\omega}.$ (18)
Here, $f(\mathcal{E},T_{0})$ is the electron distribution at $T_{0}=300$ K. The ITO permittivity (shown in Fig. 2(b)) incorporates the finite-size effect through the energy-space dipole matrix element $|D|$ (Eq. (16)), leading to a larger imaginary part compared to the bulk. Consistently, the Fourier transform of Eq. (17) reproduces the standard quasi-electrostatic solution for the internal field of a small spherical NP [65], exhibiting a relatively broad resonance centred at $\sim 180$ THz, see Fig. 2(b).
@@
The electric field of the incident pulse is assumed to have a Gaussian envelope, i.e., ${\bf E}_{\textrm{inc}}(t)=\hat{{\bf z}}E_{0}e^{-2\ln 2(t/\tau_{p})^{2}}\cos(\omega_{0}t)$, where $E_{0}$ is the peak amplitude of the incident pulse. We set the pulse duration to $\tau_{p}=5$ fs and choose a central frequency $\omega_{0}/2\pi=175$ THz (equivalently $\sim 0.723$ eV), close to the resonance (see Fig. 2(b)), corresponding to an optical cycle of approximately 6 femtoseconds. The incident pulse is assumed to be polarized in the $z$ direction without loss of generality. However, for simplicity, from now on we suppress the vectorial nature of the electric fields.
@@ III.1 Term comparison
We first plot in Fig. 3(a)-(c) the photon-e, e-e, and e-ph terms of Eq. (7a) for $E_{0}=0.6$ V/nm. One can see that, due to the dominance of near-diagonal terms in the dipole transition elements (see Fig. 1), the photon-e term exhibits distinct thermal-like features [24, 36] in the vicinity of the Fermi energy (see Fig. 3(a)-(c)). Additionally, the effect of electron collisions with other electrons is an order of magnitude weaker, and the effect of electron collisions with phonons is a further order of magnitude smaller. Fig. 3(d) further shows that throughout most of the pulse duration, the photon–e term overwhelmingly dominates over both scattering processes. Consequently, the ultrafast (few-femtosecond) evolution of the electron distribution is primarily driven by the photon–electron interaction. The e–e term becomes appreciable only toward the tail of the pulse ($t\gtrsim 15$ fs), after the peak of the optical excitation (see inset of Fig. 3(d)). In [1], we focused on the analysis of the overall dynamics and the two main competing contributions to this dominant photon-e term, namely, absorption and stimulated emission. Here, instead, we focus on aspects not discussed in [1]; specifically, after justifying the neglect of the contribution of spontaneous emission to photon-e interactions (see Appendix. A), we here dwell on the microscopic details of the (stimulated) photon-e interactions, namely, their energy and time dependence. This will include the relation of this term to the coherence dynamics and the origins of the super-linear scaling of the photon-e term with the illumination intensity.
Figure 3: (a)-(c) The photon-e (blue), e-e (orange), and e-ph (green) terms for $E_{0}=0.6$ V/nm at t=3.2t=-3.2 fs, 0.5-0.5 fs, and $2.2$ fs. (d) The time-dependence of photon-e (blue), e-e (orange), and e-ph (green) terms at $\mathcal{E}=\mathcal{E}_{F}+\hbar\omega_{0}/2$. The insets show the zoom-in for $t>10$ fs. (e) The color map of the photon-e term (Eq. (13)). The black dashed lines label the time sections of (a)-(c). The blue and orange dashed lines label the energy sections shown in (f). (f) The electron distribution dynamics at $\mathcal{E}=\mathcal{E}_{F}\pm\hbar\omega_{0}/8$ (blue and orange).
@@
Thus, specifically, Figs. 3(e) and (f) reveal time-oscillations at frequency $\sim 2\omega_{0}$ (i.e., with a period of $\sim 3$ fs) in this term. These arise from the product between the $\pm\omega_{0}$ components of the real electric field and the DM off-diagonal elements, which themselves oscillate at $\pm\omega_{0}$ as driven by the field (see Eqs. (13) and (14)). This $\sim 2\omega_{0}$ oscillation in the photon-e term, in turn, induces similar oscillations in the distribution. Simultaneously, cross-terms between the $\pm\omega$ components of $E(t)$ and the $\mp\omega$ components of the DM off-diagonal elements generate a zero-frequency component in the photon-e term, resulting in a non-vanishing, time-averaged change in the distribution. As the field intensity increases, the electron distribution oscillation stimulates higher-order oscillations in the off-diagonal density matrix at $\sim\pm 3\omega_{0}$, $\sim\pm 5\omega_{0}$, etc., via Eq. (10b). This cascade gives rise to additional photon-e and distribution oscillations at $4\omega_{0}$, $6\omega_{0}$, and beyond, see Fig. 4. These high-frequency oscillations are typically neglected in the rotating wave approximation; however, in the current context, they signify the crucial role played by the off-diagonal elements of the density matrix in the time evolution. Notably, the off-diagonal elements are mostly dominated by those near the diagonal, with magnitudes comparable to the diagonal elements. More specifically, only off-diagonal elements near the Fermi energy are non-zero at the leading edge of the pulse, while elements further below the Fermi energy become pronounced towards the trailing edge of the pulse, see Fig. 5.
@@ III.2 Excited state absorption
$f(\mathcal{E},t)-f(\mathcal{E}^{\prime},t)=f(\mathcal{E},t)(1-f(\mathcal{E}^{\prime},t))\Theta(\mathcal{E}^{\prime}-\mathcal{E})+f(\mathcal{E},t)(1-f(\mathcal{E}^{\prime},t))\Theta(\mathcal{E}-\mathcal{E}^{\prime})\\ +(1-f(\mathcal{E},t))f(\mathcal{E}^{\prime},t)\Theta(\mathcal{E}-\mathcal{E}^{\prime})+(1-f(\mathcal{E},t))f(\mathcal{E}^{\prime},t)\Theta(\mathcal{E}^{\prime}-\mathcal{E}),$ (21)
where $\Theta$ is the Heaviside step function. If we then multiply each of these four terms by the electron energy and the eDOS, and then integrate them over time (rather than on electron energy as in Eq. (19)), the resulting quantities represent the individual contributions of each process to the total energy change of the electron system, i.e.,
@@
i.e., the sum of all four contributions, $u_{\textrm{abs,in}}$, $u_{\textrm{abs,out}}$, $u_{\textrm{em,in}}$, and $u_{\textrm{em,out}}$, integrated over $\mathcal{E}$ recovers the total energy transferred via the photon–e interaction, i.e., Eq. (20). Figs. 7(b)-(e) show the contributions from absorption (in and out) and emission (in and out) for $E_{0}=0.6$ V/nm and $1.2$ V/nm. At the lower illumination level ($E_{0}=0.6$ V/nm), absorption is dominated by excitation from below the Fermi level to higher-energy states. As the field strength increases, this contribution weakens, and excitations from already excited electrons ($\mathcal{E}>\mathcal{E}_{F}$) to even higher levels become more significant. The corresponding absorption (in) process exhibits behavior similar to that of the absorption (out) counterparts. In particular, the energy-resolved contribution $u_{\textrm{abs (in)}}(\mathcal{E})$ peaks near $\mathcal{E}_{F}+\hbar\omega_{0}$. As the illumination level increases, the peak slightly decreases, while the contributions at higher energies ($\mathcal{E}\gtrsim\mathcal{E}_{F}+\hbar\omega_{0}$) become more prominent, again indicating the onset of excited state absorption processes, see Figs. 7(b) and (c). This behavior indicates that the superlinear scaling of (stimulated) absorption with incident power originates from a non-simultaneous two-photon absorption process, analogous to excited-state absorption in atomic systems. It reflects the increasing availability of high-energy absorption channels at elevated field strengths and contrasts sharply with the saturation behavior typical of two-level systems, where absorption eventually becomes sublinear due to ground-state depletion.
In contrast, the emission (out) (and (in)) contributions, when normalized by $|E_{0}|^{2}$, increase with incident power, but their energy dependence remains largely unchanged, see Figs. 7(d) and (e). This is consistent with the observed $|E_{0}|^{4}$ scaling of emission shown in Fig. 4(e) of Ref. [1]. Notably, such behavior can only be captured by explicitly tracking the time-dependent population dynamics and cannot be reproduced by phenomenological models.
@@ III.3 Local field
Figure 9: (a) The phase of the prefactor of the linear quasi-static solution for the local field inside a subwavelength dielectric sphere; the vertical dashed line represents the central frequency ($175$ THz) used in the simulations. (b) The phase difference between the local field $E_{\textrm{NP}}(t)$ (17) and the incident field for the three cases shown in Fig. 8. (c) The phase difference between the local and incident fields under weak illumination (i.e., in the linear regime) for three pulse durations: 5 fs (solid), 20 fs (dashed), and 80 fs (dash-dotted).
@@ IV Outlook
In this work, starting from the density-matrix equations formulated in momentum space, we develop an energy-space density-matrix framework for conduction electrons driven by intense single-cycle optical pulses. Applying this approach to an ITO nanosphere, we show that the nonlinear response extends well beyond simple population redistribution: coherences between conduction-band states give rise to pronounced oscillatory dynamics and repeated frequency-doubling features in the polarization. We further identify clear signatures of coherence in the absorption and permittivity dynamics. We also reveal the emergence of strong excited-state absorption which is responsible for the superlinear scaling of the absorbed energy; this behaviour is the opposite to the sublinear behavior typically associated with ground-state depletion in two-level systems, and it mediates the sublinear growth of the net absorption. Finally, we systematically characterize the impact of pulse intensity on the temporal duration, spectral broadening and shift, and phase of the local field, thereby establishing the predominantly absorptive nature of the nonlinear response induced by single-cycle excitation in TCOs.
This work paves the way to more comprehensive studies of the nonlinear optical response for different particle sizes and shapes (as, e.g., in [52]), different material parameters or TCOs (effective mass, non-parabolicity, electron density, …) different geometries (primarily, of layers, which is the geometry of choice in most experiments to date), for different pump wavelengths (e.g., for a $800$ nm, as in [19]), probe wavelength and their duration and chirp. It would also be a basis for comparison with different more computationally accessible approaches (e.g., the hydrodynamic approach, see [75]), thus, enabling further studies of complex spatio-temporal dynamics of single cycle pulses in TCOs and interplay of coherent wave mixing with permittivity changes studied therein.